### v1

# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

Abstract
Study Objectives
We investigate using Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a minimally intrusive dual-sensor wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and temperature, as well as finger photoplethysmography (PPG) and temperature.
Methods
We obtained wearable sensor recordings from 360 adults undergoing concurrent clinical polysomnography (PSG) at a tertiary care sleep lab. PSG recordings were scored according to AASM criteria. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained Mamba-based models with both convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings. Ensembling of model variants with similar architectures was performed.
Results
Our best approach, after ensembling, attains a 3-class (wake, NREM, REM) balanced accuracy of 83.50%, F1 score of 84.16%, Cohen’s of 72.68%, and a MCC score of 72.84%; a 4-class (wake, N1/N2, N3, REM) balanced accuracy of 74.64%, F1 score of 74.56%, Cohen’s of 61.63%, and MCC score of 62.04%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 64.30%, F1 score of 66.97%, Cohen’s of 53.23%, MCC score of 54.38%.
Conclusions
Deep learning models can infer major sleep stages from a wearable system without electroencephalography (EEG) and can be successfully applied to data from adults attending a tertiary care sleep clinic.
1University of Toronto Dept. of aMedicine, bMedical Biophysics, cComputer Science Toronto, ON, Canada
2Sunnybrook Research Institute aDept. of Medicine Neurology Div., bPhysical Sciences Toronto, ON, Canada
3Vector Institute, 4University Health Network Toronto, ON, Canada
5Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
[…]
andrew.lim@utoronto.ca
Keywords: Deep Learning Sleep Staging Wearable Devices
1 Introduction
1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in the diagnosis of neurodegenerative diseases [1]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment. This is followed by sleep staging, the manual categorization of time periods into different stages of sleep based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three broad stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep. EEG is the principal signal used for sleep staging [2, 3, 4].
1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may also be uncomfortable for many patients and can result in sleep that is unrepresentative of that of the home environment. Furthermore, the throughput of PSGs is limited by the time-consuming process of monitoring patients throughout the night and then visually labelling sleep epochs.
In response to these challenges, wearable devices have been developed that record some but not all of the signals recorded by conventional PSG. Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head, similar to PSG. On the other hand, EEG-less devices are less obtrusive, and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [7], Transformers [8], and graph neural networks [9]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that take in a sequence of vectors and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
Mamba [14] provides a solution to both problems as a type of state space model (SSM) [15] that combines the excellent long-range dependency modeling of SSMs with an additional time-dependent selection mechanism that allows for focusing on or ignoring particular features, as well as a hardware-friendly algorithm that provides computational efficiency. Empirically, Mamba has achieved state-of-the-art performance on supervised and unsupervised metrics using datasets of text, audio waveforms, and DNA sequences [14], while Mamba-based models have been adopted to perform automatic sleep staging using EEG and other PSG signals to high degrees of accuracy [16, 17].
![[Uncaptioned image]](extracted/6084238/figures/anne.png)
| Statistic | Range / Count | ||
| Recordings | 357 | ||
| Male/Female | 181/176 (51/49%) | ||
| Awake [%] | 43.43 | 17.15 | 11.54 95.24 |
| N1 Sleep [%] | 10.67 | 6.13 | 1.05 33.13 |
| N2 Sleep [%] | 27.18 | 10.73 | 0.58 63.59 |
| N3 Sleep [%] | 10.38 | 7.94 | 0.00 54.83 |
| REM Sleep [%] | 8.34 | 5.81 | 0.00 28.80 |
| Age [y] | 57.49 | 17.56 | 18.17 92.07 |
| Apnea-Hyponea Index (AHI) | 11.54 | 17.21 | 0.00 127.00 |
| Total Sleep Time (TST) [h] | 5.10 | 1.46 | 0.66 9.16 |
| Body Mass Index (BMI) [kg ] | 27.81 | 5.73 | 10.00 50.40 |
| Periodic Limb Movement Index (PLMI) | 15.01 | 26.99 | 0.00 169.50 |
| Sleep Efficiency [%] | 71.83 | 17.91 | 11.60 98.50 |
1.4 Contribution
This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-sensor wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature (Temp.), and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry (Acc.), and temperature. The raw signals of a typical recording are plotted in Figure 1B.
An anterior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance. We aim to develop a better population model that allows accurate sleep staging from ANNE One signals recorded from previously unseen patients.
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. Finally, this work broadens the Mamba architecture to EEG-less sleep staging, and aims to answer whether the incorporation of Mamba offers a relative performance increase similar to PSG EEG sleep staging as noticed in [16, 17].
2 Method
2.1 Dataset
ANNE wearable sensors were used to collect full-night recordings from 360 adults undergoing concurrent clinical PSG at a tertiary care clinical sleep laboratory in Sunnybrook Health Sciences Centre, Toronto, Canada. PSG recordings were recorded according to AASM practice parameters for in-labratory PSGs using a Grael PSG system (Compumedics, Victoria, Australia). The study was conducted in accordance with the latest version of the Declaration of Helsinki. All subjects provided written informed consent for the full study protocol.
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labelled by one of four technologists according to AASM standards [2]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, leaving a total of 357 patients. The characteristics of the study’s participants are outlined in Figure 1C.
The machine learning task is thus defined from the dataset as multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
2.2 Preprocessing
The ECG data were recorded at 512 Hz and preprocessed as follows. A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. Meanwhile, the Hilbert transform was applied to extract the envelope of each QRS complex, the area under which being taken as a measure of the signal to noise ratio (i.e. signal quality index (SQI)). For each peak, morphological features (e.g. amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters. Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes. QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
The PPG data were recorded at 128 Hz and preprocessed as follows. First, PPG signals were band-pass filtered between and downsampled to 100 Hz. Trains of physiological pulse waves were identified on the basis of relatively stable HR and amplitude. These high-quality pulse waves were normalized to a standard amplitude and duration, and clustered using HDBSCAN. For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster; beat-to-beat template matching was performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave. For each matching pulse wave, the timing of the foot and peak was noted, from which a measure of heart rate was obtained.
The pulse arrival time (PAT) is the delay from the R-peak of the ECG to the arrival of the pulse wave in the PPG. To compute PAT from the ANNE data, the time-series of valid ECG peaks and PPG feet were matched to each other by empirically considering all possible lags within the range msec and selecting that which successfully matched the greatest number of ECG peaks and PPG feet. Afterwards, for each ECG R-peak and its corresponding PPG foot, the PAT of that heartbeat was computed, generating a beat-to-beat time series of PAT.
The chest XYZ accelerometry data were recorded at 210 Hz and low-pass filtered, followed by the computation of the euclidean norm minus one (ENMO), theta (body pitch), phi (body roll) and the z-angle (the deviation from vertical of a vector perpendicular to the face of the sensor).
2.3 Model
Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3. The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs. The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs. 70K). Since Mamba is designed for pattern recognition across long contexts, it could potentially interact with signals more effectively than the CRNN if the signals are given directly on a finer time-resolution without windowing the data.
The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows, the window size (i.e. 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features. The model comprises:
1)
A convolutional component, consisting of 1D residual blocks that sequentially downsize the time-series and frequency datas’ window size axes and flatten them into an embedding vector per window. The time-series embedding, frequency embedding and scalar data are then concatenated across the feature axis to form an embedding vector per window.
2)
A recurrent component, consisting of a bidirectional Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class. The bidirectional Mamba block (Appendix A) is the concatenation of two Mamba blocks that process epochs in opposite directions.
The RNN architecture (Figure 2B) processes a single 2D input channel with axes for the number of -second windows and number of features, where is an arbitrary value dividing 30. It comprises:
1)
An element-wise MLP with hidden layer dimensions that downsizes the data to dimensions, where the choices of and are arbitrary.
2)
A bidirectional Mamba block that processes windows sequentially.
3)
An element-wise MLP that downsizes the data into logits, which, during inference-time, can be grouped and averaged to form logits per 30-second window to comply with AASM standards.
Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches. The engineered scalar features distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix B). Thus, empirical selections of these features (Table 1) were introduced alongside the time-series signals for both the CRNN and RNN.
A series of transformations were applied to further adjust the preprocessed data to the individual architectures. For the CRNN, all signals were resampled to 25 Hz and grouped to 30-second windows, resulting in data points per window. Then, depending on feature type (see Table 1),
For time-series features, the signal was left alone.
For frequency features, the log-magnitude spectrum was taken of the detrended signal over the 30-second axis.
For scalar features, the signal was reduced over the 30-second axis. Metadata, in the form of patient age and sex (male = 0, female = 1), was incorporated into the scalar features by repeating the constants times throughout the recording.
For the RNN, the accelerometry, ECG, and PPG signals were detrended, with the accelerometry signals further band-pass filtered over Hz. Log-magnitude spectrograms were taken with a window size of 100 points and no overlap, then finally concatenated with limb/chest temperature, which was resampled to 1 Hz, and sleep stage label, of which each point was repeated 30 times to 1 Hz.
RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO) Z* Entropy, PPG Chest Temp. Z-Angle Z-Angle* Entropy, PPG* Limb Temp. Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle
2.4 Training
The training, validation, and testing sets were randomly split in a size ratio to sizes . During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, which mostly consisted of wake periods at the start and end of the recording with sleep in the middle. Cross-entropy loss was weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C. The model from the iteration that maximized balanced validation accuracy was chosen as the final model. Both CRNN and RNN models were trained with the Adam optimizer [32].
The CRNN model was trained with single batches. Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation. The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
Three models of varying sleep stage resolution were built: a 3-class, 4-class, and 5-class sleep staging model. The 5-class model has classes for wake, N1, N2, N3, and REM sleep, while the 3-class model groups N1, N2, and N3 sleep into non-REM sleep. The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [4].
2.5 Ensembling
For each of the classes inside a -class ensemble model, a binary model was trained to predict the presence or absence of that class. The logits of these binary models were stacked with logits from the regular -class models to form the inputs to a suite of decision tree-based architectures including AdaBoost [33], random forests [34], extra-tree classifiers [35], histogram gradient boosting [36], and XGBoost [37]. The model among the suite that maximized balanced validation accuracy was selected as the final ensembler model. A visual representation of the ensembling pipeline is presented in Figure 2C.
As decision tree-based architectures do not output class-wise logits, epoch-wise predictions on the RNN model were generated by aggregating 30 one-second predictions and reducing with the mode.
3 Results
Results are compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models. Overall performance is evaluated by aggregating all test set recordings and computing the balanced accuracy (B.Acc.), precision (Prec.), recall (Rec.), F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC). We additionally conduct analyses of averaged per-recording metrics, biases in the amount of each sleep stage, and differences in performance across age, PLMI, AHI, BMI, TST, and sleep efficiency. Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy (Acc.) is presented in lieu of balanced accuracy for those cases.
3.1 Model Evaluation
Typical-case recording predictions for all 5-class models are shown in Figure 4 along with predicted point-wise output class probabilities for models that can produce them. A Uniform Manifold Approximation and Projection (UMAP) [38] of an intermediate CRNN output is also provided in Appendix E, which shows the separation of sleep stages as clusters in the learned representation space.
Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. No significant deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows, of which the latter is presented in this paper. The RNN ensemble model yields the best metrics and dominates most class-wise metrics (Appendix D). Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count. Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
Lastly, in a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model’s performance is significantly boosted as denoted in the confusion matrices and metrics of Figure 6.
![[Uncaptioned image]](extracted/6084238/figures/ensemble_outputs.png)
[…]
![[Uncaptioned image]](extracted/6084238/figures/healthy-subset.png)
Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.26 76.85 70.07 71.55 60.48 61.60 4 79.39 78.82 77.00 76.89 67.23 67.92 3 86.32 87.63 86.70 86.88 77.87 78.14
3.2 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs. PSG actual values for a number of summary metrics including time asleep, % N1 sleep,% N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating N2 sleep and slightly overestimating N1 sleep and wake.
3.3 Performance by Clinical Group
Averaged recording-wise metrics are cataloged in Table 2 for the 5-class ensemble CRNN and RNN, wherein linear regressions were performed to fit various recording metrics to accuracy. Performance of the RNN ensemble is robust to differences in sex, periodic limb movement index, body mass index, total sleep time and sleep efficiency, but shows slightly worse performance with greater age and higher AHI. The CRNN model is robust to differences in age, sex, PLMI, and BMI but shows slightly worse performance with higher AHI, shorter TST, and lower sleep efficiency. Notably, even where performance varies by clinical factors, the percentage of explained change in performance is nevertheless weak for all metrics (). There is no significant difference in accuracy between the raters in the testing set (RNN , CRNN , one-way F-test).
| RNN | CRNN | ||||||
| Metric (%) | |||||||
| Accuracy | 65.09 | 9.84 | 62.58 | 12.50 | |||
| F1 | 67.69 | 9.92 | 63.81 | 14.17 | |||
| Cohen’s | 50.39 | 14.91 | 46.33 | 16.13 | |||
| Matthews Correlation Coefficient | 52.61 | 13.94 | 48.77 | 15.91 | |||
| Sleep Statistic | |||||||
| Age | 0.0827 | 0.0350 | 0.0346 | 0.1779 | |||
| Sex | 0.0689 | 0.0552 | 0.024 | 0.2630 | |||
| Periodic Limb Movement Index (PLMI) | 0.0120 | 0.4300 | 0.0012 | 0.8061 | |||
| Apnea-Hypopnea Index (AHI) | 0.1525 | 0.0350 | 0.1622 | 0.0025 | |||
| Body Mass Index (BMI) | 0.0164 | 0.3559 | 0.0000 | 0.9824 | |||
| Total Sleep Time (TST) | 0.0000 | 0.9921 | 0.0853 | 0.0322 | |||
| Sleep Efficiency | 0.0045 | 0.6316 | 0.1503 | 0.0038 | |||
4 Discussion
4.1 Key Findings
This work has successfully developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages. The models demonstrated good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 3 & 4) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity. In contrast, we studied a much older (mean age 57.49) participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6). Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality may be related in part to its use of multiple streams of data such that sensors returning poor quality data are compensated for in the model by those returning good quality data.
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the person-specific models generated in this paper.
[…]
4.3 Model Performance
Both model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models). Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
In spite of a simpler feature set and smaller parameter count, the RNNs outperform the CRNNs in virtually every metric. It is possible that the independent processing of segmented windows by the CNN component interfered with the wider CRNN’s ability to discern changes at neighbouring window boundaries. As well, during epochs where sleep is transitioning between classes, the provided label becomes less reliable; the CRNN attempts to predict a behaviourally inconsistent epoch while it is likely the RNN more flexibly captures this ambiguity by means of multiple differing fine-resolution class predictions throughout the 30 seconds.
No statistically significant variations in performance based on PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts. The significant () association discovered between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in more time in N1 sleep, one of the most-difficult stages to separate; even so, the association was modest (). Model performance also correlates between age and accuracy (), though this may in part be explained by the weak agreement between AHI and age ().
4.4 Limitations, Strengths, and Future Directions
A few methodological limitations are worth considering. 1) The differing nature of the way our CRNN and RNN models handle preprocessing and predictions (notably, the 30-second window employed by the CRNN as opposed to the 1-second window used in the RNN) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance. 2) As neural networks, CRNNs and RNNs provide limited insight into the decoding of sleep stages from the given sensor data. Further analyses targeted toward interpreting the neural networks may be explored. 3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake. It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative.
This study also had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the combination of Mamba with an RNN architecture allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
5 Conclusion
This work introduces CRNN and RNN models to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings, with the RNN ensemble model topping the others in virtually every performance metric. The model meets or surpasses metrics in related literature, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics. Moreover, the model’s capability was achieved without prescreening and excluding large numbers of patients or epochs for reasons of signal quality. Compared to PSGs, the ANNE One sensor is far less intrusive of a device and procedure, which is significant for expanding the accessibility of sleep measurement for older and remote-living adults at greater risk of sleep-related illness. Our model provides a scalable, accurate minimally-intrusive approach to the ambulatory assessment of sleep staging without the need for EEG.
Acknowledgments
This research is supported by the Centre for Aging and Brain Health Innovation, Canadian Institutes of Health Research, and National Institute on Aging.
This research is not funded by Sibel Health, the manufacturer of the ANNE One device.
Special thanks to Prof. Jimmy Ba from the Department of Computer Science; Prof. Paul Kushner and Prof. Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
Appendix A: Bidirectional Mamba Block
Appendix B: Scalar and Frequency Features
Appendix C: Explorations in Feature Selection
A previous iteration of the model used a pure CRNN (GRU) approach, and the approach encompassed two additional strategies, which are utilized within the paper without extensive justification.
The engineered features are a suite of additional features across the time, frequency, and scalar domains used alongside the raw signals as extracted from the ANNE sensors.
Time Domain Frequency Domain Scalar Domain PPG, ECG Heart Rates PPG Freq. Spectra PPG Freq. Spectra Skew PPG, ECG Signal Quality Index ECG Freq. Spectra PPG Time Entropy PPG, ECG Amplitude x-accel. Freq. Spectra PPG Freq. Spectra Entropy Pulse Arrival Time (PAT) y-accel. Freq. Spectra | Heart Rate| Accel. ENMO z-accel. Freq. Spectra Accel. ENMO Kurtosis SpO2 Accel. z-angle Freq. Spectra Entropy Metadata (Gender, Age)
In addition, the auxiliary training objective was introduced in the form of the transition state. These states are derived directly from the target labels. With this, the model attempts to predict whether there will be a change in sleep class from this epoch to the next. The rationale is that sleep stages in reality do not always change at the end of 30-second windows, and many epochs may have data that is transitioning between two separate sleep stages, which would be difficult for the model as they may not resemble other members of their labelled class and instead as a mixture of both sleep stages. By training the model to also predict correct transitions, it should assist the model in determining which epochs may be occurring during a transition between stages, and thus make it more robust in general.
The application of the transition state as an auxiliary objective was attempted in the Mamba approaches in the main paper; however, it did not yield a noticeable increase in performance, and thus was not included in the final models.
The results of the approach are detailed in the graph below, with results indicating that the two approaches mentioned positively impacted results.
3-class Sleep Staging 2-class Sleep Detection F1 Precision Accuracy F1 Precision Accuracy Wake Time Series Only Baseline 0.681 0.723 0.644 0.740 0.649 0.860 Engineered Features 0.743 0.723 0.764 0.742 0.703 0.786 Engineered Features & Auxiliary Objective 0.727 0.757 0.700 0.781 0.701 0.880 Ensemble 0.732 0.789 0.683 NREM Sleep Time Series Only Baseline 0.534 0.720 0.425 0.793 0.892 0.714 Engineered Features 0.507 0.754 0.381 0.826 0.859 0.826 Engineered Features & Auxiliary Objective 0.619 0.744 0.530 0.834 0.912 0.769 Ensemble 0.702 0.767 0.647 REM Time Series Only Baseline 0.401 0.265 0.827 Engineered Features 0.428 0.287 0.844 Engineered Features & Auxiliary Objective 0.447 0.314 0.778 Ensemble 0.517 0.381 0.679 Macro Time Series Only Baseline 0.539 0.632 0.555 0.767 0.771 0.770 Engineered Features 0.559 0.663 0.581 0.783 0.784 0.792 Engineered Features & Auxiliary Objective 0.600 0.598 0.624 0.807 0.807 0.811 Ensemble 0.651 0.646 0.679
Appendix D: Class-Wise Evaluation Metrics
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC Wake Wake Wake CRNN 83.52 84.10 84.13 84.07 67.46 67.53 84.59 84.80 84.75 84.77 69.01 69.02 84.57 84.70 84.45 84.50 68.57 68.70 RNN 86.00 88.85 87.59 87.29 74.01 75.63 88.00 89.57 89.08 88.94 77.36 78.05 88.67 89.26 89.23 89.18 77.90 78.01 CRNN Ens. 85.63 86.90 86.64 86.49 72.36 72.83 86.96 87.40 87.41 87.38 74.24 74.28 88.05 88.52 88.52 88.49 76.49 76.55 RNN Ens. 86.68 88.63 87.95 87.75 74.93 75.88 88.49 89.14 89.10 89.04 77.61 77.74 89.03 89.41 89.42 89.39 78.36 78.39 N1 Sleep N1/N2 Sleep Non-REM Sleep CRNN 66.02 86.62 79.98 82.65 22.47 24.23 62.09 67.98 68.51 65.69 26.58 29.20 79.70 80.20 79.98 79.88 59.68 60.00 RNN 67.77 87.06 77.97 81.43 22.61 25.40 68.42 71.65 72.13 71.34 38.39 39.00 81.68 82.36 82.00 81.89 63.72 64.18 CRNN Ens. 68.82 87.42 80.47 83.18 25.87 28.12 71.45 74.71 75.07 74.32 44.79 45.56 82.29 82.67 82.52 82.46 64.83 65.05 RNN Ens. 68.96 87.42 79.35 82.44 24.96 27.61 73.52 76.06 76.40 75.96 48.45 48.83 85.15 85.35 85.30 85.28 70.47 70.56 N2 Sleep CRNN 65.60 74.43 76.10 74.58 34.46 35.48 RNN 64.82 76.06 77.46 74.93 34.63 37.49 CRNN Ens. 65.29 75.54 77.14 75.00 35.03 37.10 RNN Ens. 68.12 77.33 78.59 77.02 40.56 42.13 N3 Sleep N3 Sleep CRNN 73.95 90.34 89.41 89.83 44.03 44.26 75.29 89.62 81.09 84.14 31.74 35.77 RNN 82.25 91.85 87.58 89.08 47.45 50.21 82.22 91.58 84.59 86.94 42.11 46.62 CRNN Ens. 77.66 91.19 89.59 90.25 48.04 48.69 78.35 91.05 88.33 89.39 45.80 47.15 RNN Ens. 81.20 92.14 90.30 91.01 52.80 53.78 82.18 92.12 89.32 90.35 51.18 52.91 REM Sleep REM Sleep REM Sleep CRNN 72.17 91.05 91.33 91.18 45.86 45.90 73.16 90.10 87.77 88.75 38.54 39.38 78.98 92.24 91.29 91.70 52.43 52.79 RNN 82.42 93.23 92.29 92.67 58.19 58.66 82.85 93.17 91.90 92.40 57.33 58.04 87.46 93.49 88.87 90.37 52.70 56.55 CRNN Ens. 81.70 92.69 91.06 91.70 54.01 54.92 83.06 92.82 90.54 91.38 53.64 55.13 83.47 92.73 89.69 90.78 51.80 53.90 RNN Ens. 84.79 94.20 93.63 93.86 64.41 64.68 87.16 94.40 93.22 93.65 64.47 65.33 88.17 94.49 92.94 93.48 61.16 65.42
Appendix E: UMAP of CRNN Model Learned Embeddings